### v1

# Broadband directional thermal emission with anisothermal microsources

Abstract
Thermal emission is naturally spatially incoherent and lacks directionality. Here, we demonstrate that by precisely controlling the spatial temperature distribution within a solid, directional thermal emission can be achieved across a broad spectral range. These anisothermal sources open new avenues for manipulating radiative heat flux at the microscale and hold promise for applications in thermal management and energy conversion, enabling more efficient and targeted thermal control.
pacs:
44.40.+a, 78.20.N-, 03.50.De, 66.70.-f
Thermally generated light is a fundamental phenomenon resulting from the agitation of partial charges in matter due to the current fluctuations induced by the local temperature. Its control is crucial for numerous technologies like energy conversion, imaging or sensing. Traditional thermal emitters are incoherent Planck ; Kirchoff and lack directionality due to the random nature of emission process. Various strategies, such as surface-polariton gratings Hesketh86 ; Hesketh88 ; Kreiter ; Greffet , metasurfaces Wang , cavities Kollyukh ; PBA_JOSA ; Celanovic , and photonic crystals Lee ; Battula have demonstrated angular control of thermal emission, but this control is limited to narrow bandwidths. Achieving broadband thermal emission confined to specific directions is a much more complex problem since it presents several fundamental challenges. Indeed, thermal emission spans a wide range of wavelengths and controlling the directionality across this broad spectrum is inherently difficult since matter structuring is generally well adapted to the control of specific wavelengths. However, recently several mechanisms have been proposed to achieve a broadband directional thermal emission. Among these is the perfect absorption at the Brewster angle in hyperbolic materials of type II over a large spectral domain Barbillon . However these sources emit only p-polarized light, which intrinsically limits their flux. A second mechanism is the opening of angular band gaps Hamam ; Shen with stacks of anisotropic photonic crystals. In such systems light can only propagate in specific directions of space, all other directions being forbidden due to destructive interferences. Finally, gradient epsilon-near-zero materials have been proposed Xu to achieve broadband directional control. These media support leaky p-polarized Berreman modes able to couple to free-space modes at specific angles across a wide frequency range. However, the practical deployment of these directive sources is hindered by challenges such as fabrication complexity and material losses. Moreover, they do not allow an active change of the direction of emission once the source is designed. In the present Letter, we introduce the concept of anisothermal sources, which are sources with spatially variable temperatures profiles. By analyzing their emission, we demonstrate that directivity of radiative heat flux can be achieved through a precise control of their temperatures, showing in this way the tunability potential of these thermal antennas.
To start, we consider a finite-size object of volume $\mathcal{V}$ and of frequency-dependent permittivity $\epsilon(\omega)$ immersed in a thermal bath at temperature $T_{\mathrm{b}}$. We also assume that the temperature distribution $T(\mathbf{r})$ within this object is non-uniform. The thermal light radiated by this object toward its surrounding environment can be calculated using the fluctuational-electrodynamics theoretical framework Rytov . This calculation can be done easily by using the discrete-dipole approximation method pba2011 ; Edalatpour ; Cuevas which consists in discretizing the solid into a finite number $N$ of small polarizable fluctuating dipoles arranged on a grid. Each dipole represents a small volume element of the material and responds to the local electromagnetic field. Those dipoles interact with each other and with the external bath. The local electric and magnetic fields $\mathbf{E}$ and $\mathbf{H}$ in the surrounding environment are linearly related to fluctuating dipolar moments $\mathbf{p}_{j}^{\mathrm{(fl)}}$ and to the fields $\mathbf{E}_{\mathrm{b}}$ and $\mathbf{H}_{\mathrm{b}}$ radiated by the bath as follows pba2011 ; messina2013 ; biehs1
$\begin{split}\mathbf{E}(\mathbf{r},\omega)&=\mathds{B}^{\mathrm{(E)}}(\mathbf{r},\omega)\mathbf{E}_{\mathrm{b}}(\mathbf{r},\omega),\\ &\,+\mu_{0}\omega^{2}\sum_{j=1}^{N}\mathds{G}^{\mathrm{(EE)}}(\mathbf{r},\mathbf{r}_{j},\omega)\mathbf{p}_{j}^{\mathrm{(fl)}}(\omega),\end{split}$ (1)
and
$\begin{split}\mathbf{H}(\mathbf{r},\omega)&=\mathds{B}^{\mathrm{(H)}}(\mathbf{r},\omega)\mathbf{H}_{\mathrm{b}}(\mathbf{r},\omega)\\ &\,+\mu_{0}\omega^{2}\sum_{j=1}^{N}\mathds{G}^{\mathrm{(HE)}}(\mathbf{r},\mathbf{r}_{j},\omega)\mathbf{p}_{j}^{\mathrm{(fl)}}(\omega).\end{split}$ (2)
Here, $\mathds{G}^{\mathrm{(EE)}}$ and $\mathds{G}^{\mathrm{(HE)}}$, representing the full electric and magnetic dyadic Green tensors at frequency $\omega$, as well as the matrices $\mathds{B}^{\mathrm{(E)}}$ and $\mathds{B}^{\mathrm{(H)}}$ take into account all many-body interactions. These matrices can be deduced by solving a self-consistent linear system as detailed in Ref. messina2013 .
Figure 1: (a) Polar plot of Poynting flux (target) through an oriented surface of normal $\mathbf{n}=(\sin\theta,\cos\theta)$ at $z=50\,\mu$m above a dipolar chain made with 500 SiC nanoparticles $R=20\,$nm separated by a distance (center to center) $d=4R$. (b) Temperature profile calculated along the chain to achieve the target thermal emission. (c) Spatial distribution of the power required along a SiC cylinder of radius $R=20\,$nm to maintain its temperature at the value plotted in (b) when the bath is at $T_{\mathrm{b}}=300\,$K. The value $P_{\mathrm{tot}}$ shown in the inset corresponds to the total power needed to control the temperature profile and the thermal conductivity of SiC is $\lambda=35\,\mathrm{W}\cdot\mathrm{m}^{-1}\cdot\mathrm{K}^{-1}$.
Figure 2: Same results as in Fig. 1 with tilted emission patterns. (a) Targets (b) Temperatures profiles (c) Power distribution.
Figure 3: Comparison of the Poynting flux radiated by the same dipolar chains as in Figs. 1 and 2 between the linear approximation (dashed lines) and the exact (nonlinear) model (solid lines) for the narrow emission pattern (a) and the tilted one (b).
The heat flux radiated by the solid through an oriented surface can be calculated from the statistical averaging of the spectral Poynting vector
$\langle\mathbf{S}(\mathbf{r},\omega)\rangle=2\,\mathrm{Re}\langle\mathbf{E}(\mathbf{r},\omega)\times\mathbf{H^{*}}(\mathbf{r},\omega)\rangle.$ (3)
From Eqs. (1) and (2) and using the fluctuation dissipation theorem messina2013 ; Callen
$\begin{split}&\langle p^{\mathrm{(fl)}}_{i,\alpha}(\omega)p^{\mathrm{(fl)*}}_{j,\beta}(\omega^{\prime})\rangle=2\hbar\epsilon_{0}\chi_{i}(\omega)\Theta(T_{i},\omega)\delta_{ij}\delta_{\alpha\beta}2\pi\delta(\omega-\omega^{\prime}),\\ &\langle E_{\mathrm{b},\alpha}(\mathbf{r},\omega)H^{*}_{\mathrm{b},\beta}(\mathbf{r}^{\prime},\omega^{\prime})\rangle=2i\hbar\mu_{0}\omega^{2}\Theta(T_{\mathrm{b}},\omega)\\ &\qquad\times\mathrm{Re}[\mathds{G}^{\mathrm{(HE)}}_{\beta\alpha}(\mathbf{r},\mathbf{r}^{\prime},\omega)]2\pi\delta(\omega-\omega^{\prime}),\end{split}$ (4)
where $\chi_{i}=\mathrm{Im}(\alpha_{i})-|\alpha_{i}|^{2}\omega^{3}/6\pi c^{3}$, $\alpha_{i}$ being the polarizability associated to the $i$-th dipole, and $\Theta(T,\omega)={\hbar\omega}/[\exp(\hbar\omega/k_{B}T)-1]$ is the mean energy of a harmonic oscillator at temperature $T$, it is straigforward to show that each component of this vector can be recast into
$\langle S_{\zeta}(\mathbf{r},\omega)\rangle=\sum_{i=1}^{N}a_{\zeta}(\mathbf{r},\mathbf{r}_{i},\omega)\bigl{[}\Theta(T(\mathbf{r}_{i}),\omega)-\Theta(T_{\mathrm{b}},\omega)\bigr{]},$ (5)
where
$\begin{split}a_{\zeta}(\mathbf{r},\mathbf{r_{i}},\omega)&=\epsilon_{\zeta\gamma\beta}\frac{4\hbar\omega^{4}}{\epsilon_{0}c^{4}}\chi_{i}(\omega)\\ &\,\times\mathrm{Re}[\mathds{G}^{(\mathrm{EE})}_{\gamma\eta}(\mathbf{r},\mathbf{r}_{i},\omega)\mathds{G}^{\mathrm{(HE)}*}_{\beta\eta}(\mathbf{r},\mathbf{r}_{i},\omega)],\end{split}$ (6)
in which a summation is intended over the repeated indices. Eq. (5) established a close relation between the temperature profile within the solid and the spectral components of Poynting vector which it radiates in its surrounding. Assuming a weak temperature variation within the system close to the environmental temperature $T_{\mathrm{b}}$, the full Poynting vector $\langle\mathbf{S}(\mathbf{r})\rangle=\int\frac{d\omega}{2\pi}\langle\mathbf{S}(\mathbf{r},\omega)\rangle$ can be recast in the following form
$\langle\mathbf{S}(\mathbf{r})\rangle=\sum_{i=1}^{N}\int\frac{d\omega}{2\pi}\,\mathbf{a}(\mathbf{r},\mathbf{r}_{i},\omega)\frac{d\Theta(T_{\mathrm{b}},\omega)}{dT}[T(\mathbf{r_{i}})-T_{\mathrm{b}}],$ (7)
which linearly relates the Poynting vector radiated by the system to its temperature profile. This expression is the main result of this work and it can be used in inverse design to determine the precise temperature profile required to achieve a desired emission pattern. In Figs. 1 and 2 we illustrate the potential of inverse design to control the heat flux radiated by a simple dipolar chain made of silicon carbide (SiC) nanoparticles, whose optical properties are described by a Drude-Lorentz model Palik . In these examples, our objective (target) is the following Lorentzian heat flux crossing the oriented surfaces of normal unit vector $\mathbf{n}$ at a distance $z$ from the structure
$\mathbf{n}\cdot\langle\mathbf{S}_{\mathrm{targ}}\rangle=\frac{S_{0}}{1+(\frac{x-x_{0}}{\Gamma})^{2}}.$ (8)
Here $\Gamma$ and $S_{0}$ denote the half-width at half maximum (HWHM) of the Lorentzian and the maximum flux while $x_{0}$ sets the direction $\theta_{0}$ of emitted flux ($x_{0}=z\tan\theta_{0}$). The inverse problem is solved using a Tikhonov’s regularization method Tikhonov to identify the temperature profiles necessary to reach a Lorentzian flux. The solution of this problem is plotted in Fig. 1 in the case of normal emission (i.e. $x_{0}=0$) and in Fig. 2 for an oblique one, for various HWHM. As expected, the temperature profile, solution of this inverse problem, is symmetric with respect to the center of the chain for a normal emission pattern [Fig. 1(b)] while it is no longer symmetric for a tilted emission [Fig. 2(b)]. In Fig. 1 we see that, to reduce the width of emission pattern and, therefore, focus heat in a narrower lobe, the temperature contrast between the center of the chain and its edges must be increased. However, when the local temperature variation within the system becomes too significant, the solution to the problem is no longer compatible with the linear approach we have adopted. This sets fundamental limits on the control of heat flux by anisotherm sources (at least within the linear inverse problem). In Fig. 3 we show the discrepancy we obtain between the emission patterns calculated in the linear approximation of the Poynting vector in Eq. (7) and the one obtained from its full expression in Eq. (5). These curves confirm that the linear approach is sufficient if the target is reached with weakly varying temperature profiles.
Figure 4: Poynting flux through an oriented surface of normal unit vector $\mathbf{n}$ defined as in Figs. 1 and 2 at $z=50\,\mu$m above a dipolar disk of radius $l=10\,\mu$m made of 511 SiC nanoparticles having radius $R=100\,$nm distributed on a hexagonal network of side $d=8R$ for (a) a uniform temperature profile at $T=301\,$K and (b) an inhomegenous temperature profile (c) calculated by inverse design to achieve the tilted emission shown in (b). The red circle plotted in (a) and (b) delimits the countour of the system. The temperature of external bath is $T_{\mathrm{b}}=300\,$K.
To reach the target emission, the temperature profile within the solid must be carefully controlled, requiring localized heating or cooling, as shown in Figs. 1(b) and 2(b). The power density $p_{\mathrm{dis}}(\mathbf{r})$ needed to achieve such a control can be calculated from the energy balance equation
$p_{\mathrm{dis}}(\mathbf{r})=\nabla\cdot[\mathbf{S}(T(\mathbf{r}))-\lambda\nabla T(\mathbf{r})],$ (9)
where $\lambda$ sets the thermal conductivity of solid, assuming here a diffusive heat-transport regime. The power distribution $P_{\mathrm{dis}}=4/3\pi R^{3}p_{\mathrm{dis}}$ injected locally inside the system is plotted in Figs. 1(c) and 2(c) for SiC cylindrical anisothermal SiC sources in the cases of a normal and tilted emission, respectively. It is important to note that this distribution is more complex than the temperature profile itself, due to the influence of many-body interactions within the system and edge effects. Additionally, we observe that the power injected or extracted at the system’s edges dominates over that in the interior. This could present a practical challenge for temperature control, which should be addressed in future studies. Furthermore, strong thermal conduction may homogenize temperature gradients, thereby diminishing spatial control over emission directionality. In contrast, weak conduction helps maintain distinct thermal contrasts, preserving directional effects.
Finally, we illustrate the possibility to achieve a heat flux control using a two-dimensional system composed of a regular array of SiC nanoparticles arranged on a disk. As shown in Fig. 4, even a small local temperature disturbance around the equilibrium temperature $T_{\rm eq}=T_{\rm b}$ can significantly alter the direction of the emitted flux [Fig. 4(b)] compared to a system at uniform temperature [$T=301\,$K in Fig. 4(a)]. Hence, a local temperature variation of $\Delta T\approx 0.3\,$K can change the direction of emission by $\Delta\theta\approx 15$ deg.
In conclusion, we have theoretically shown that controlling spatial temperature distributions within a solid can enable broadband, directional thermal emission. This approach utilizes anisothermal sources to overcome the inherent spatial incoherence of thermal radiation, providing a method for tailoring radiative heat flux at the microscale. By strategically designing temperature gradients, it is possible to achieve directional emission across a wide spectral range, without the need for complex nanophotonic structures. Notably, the directionality of thermal emission can be easily tuned by adjusting the temperature profile. Our findings open up new possibilities for advanced thermal management and energy conversion applications, where precise control of thermal radiation could improve efficiency and enable targeted heat transfer. More broadly, this work emphasizes the potential of anisothermal emitters to expand the design space for radiative energy systems by integrating microscale thermal gradients to control far-field emission.
Acknowledgements.
This work was supported by the French Agence Nationale de la Recherche (ANR), under grant ANR-21- CE30-0030 (NBODHEAT). F. H. acknowledges financial support by the Walter Benjamin Program of the Deutsche Forschungsgemeinschaft (eng. German Research Foundation) under project number 519479175.